flying to kampala with a tailwind a plane averaged 158 km/h. on the return trip the plane only averaged 112 km/h while flying back into the same wind. find the speed of the wind and the speed of the plane in still air.
Answer by stanbon(75887)

flying to kampala with a tailwind a plane averaged 158 km/h. on the return trip the plane only averaged 112 km/h while flying back into the same wind. find the speed of the wind and the speed of the plane in still air.
--------------------------------
Let plane speed be "p".
Let wind speed be "w".
---------
Equations:
p + w = 158
p - w = 112
----
2p = 270
p = 135 mph (plane speed)
---
Solve for "w":
p + w = 158
135 + w = 158
w = 23 mph (wind speed)
===============================
Cheers,
Stan H.
=============